Eaton Tool Company has fixed costs of $210,600, sells its units for $58, and has variable costs of $32 per unit.

a. Compute the break-even point.
  



b. Ms. Eaton comes up with a new plan to cut fixed costs to $160,000. However, more labor will now be required, which will increase variable costs per unit to $35. The sales price will remain at $58. What is the new break-even point? (Round your answer to the nearest whole number.)
  



c. Under the new plan, what is likely to happen to profitability at very high volume levels (compared to the old plan)?
  

  • Profitability will be less

  • Profitability will be more
Answer #1
a.
Contribution margin per unit = Sales price per unit - Variable cost per unit = 58 - 32 26
Break-even point = Fixed costs / Contribution margin per unit = 210600 / 26 8100 units
b.
Contribution margin per unit = Sales price per unit - Variable cost per unit = 58 - 35 23
Break-even point = Fixed costs / Contribution margin per unit = 160000 / 23 6957 units
c.
Answer : Profitability will be less
Explanation : At the very high volume levels, the new plan will earn lower contribution margin as compared to the old plan and the benefit of lower fixed costs of new plan will not help at the very high volume level and hence the profitability will be less.